# Scaling Laws for Predicting Downstream Performance in LLMs

### v1

@@ Abstract
Precise estimation of downstream performance in large language models (LLMs) prior to training is essential for guiding their development process. Scaling laws analysis utilizes the statistics of a series of significantly smaller sampling language models (LMs) to predict the performance of the target LLM. For downstream performance prediction, the critical challenge lies in the emergent abilities in LLMs that occur beyond task-specific computational thresholds. In this work, we focus on the pre-training loss as a more computation-efficient metric for performance estimation. Our two-stage approach consists of first estimating a function that maps computational resources (e.g., FLOPs) to the pre-training Loss using a series of sampling models, followed by mapping the pre-training loss to downstream task Performance after the critical “emergent phase”. In preliminary experiments, this FLP solution accurately predicts the performance of LLMs with 7B and 13B parameters using a series of sampling LMs up to 3B, achieving error margins of 5% and 10%, respectively, and significantly outperforming the FLOPs-to-Performance approach. This motivates FLP-M, a fundamental approach for performance prediction that addresses the practical need to integrate datasets from multiple sources during pre-training, specifically blending general corpora with code data to accurately represent the common necessity. FLP-M extends the power law analytical function to predict domain-specific pre-training loss based on FLOPs across data sources, and employs a two-layer neural network to model the non-linear relationship between multiple domain-specific loss and downstream performance. By utilizing a 3B LLM trained on a specific ratio and a series of smaller sampling LMs, FLP-M can effectively forecast the performance of 3B and 7B LLMs across various data mixtures for most benchmarks within 10% error margins.
@@ 1 Introduction
Previous efforts focus on predicting the target LLM’s pre-training loss and establish a power-law relation between the computational resource expended (e.g., floating-point operations per second (FLOPs)) and the final loss achieved (Kaplan et al., 2020; Muennighoff et al., 2024; Henighan et al., 2020). Further, we aim to predict the downstream performance in LLMs to more accurately reflect the primary concerns regarding their capabilities. The critical challenge is the emergent abilities in LLMs, which states that LLMs only exceed random performance when the FLOPs expended during training surpass task-specific thresholds (Wei et al., 2022a). Supposing a task threshold of $F_{c}$, typical methods require training $N$ LMs, expending total FLOPs $F_{t}=\sum_{i=1}^{N}\text{FLOPs}_{i}>N\times F_{c}$, to obtain $N$ effective data points, thereby necessitating significant computational resources. Fig. 1 demonstrates that the sampling LMs require more than $5\times 10^{20}$ FLOPs to perform better than random on most benchmarks, with only three data points available to fit the predictive curve across these benchmarks. Hu et al. (2023) address this challenge by significantly increasing the sampling times to compute the PassUntil of a task, basically increasing the “metric resolution” to enable the abilities to emerge earlier (i.e., reducing $F_{c}$). However, this approach faces challenges in translating the PassUntil back to the original task metric of concerns and requires huge amounts of FLOPs spent on sampling.
@@
In this work, we target the actual task performance prediction based on two intuitions: (1) Predicting the target pre-training loss is easier and achievable since there is no “emergent phase” in the pre-training loss, as extensively verified in Kaplan et al. (2020); Hoffmann et al. (2022); (2) There is an observed correlation between the pre-training loss and the downstream task performance after the “emergent point” (i.e., the pre-training loss goes below a critical threshold) (Du et al., 2024; Huang et al., 2024). Rethinking previous practice, training LM $i$ to convergence requires expending FLOPsi for obtaining a single data point. In contrast, our approach stems from a crucial insight: Collecting (pre-training loss, performance) data points at intermediate checkpoints prevents the need for fully training LMs to convergence, thereby enhancing sample efficiency. Essentially, we can collect a huge amount of useful data between the initial point of above-random performance in LMs and their convergence point for performance prediction.
@@
Thus, our approach consists of two sequential stages: (1) $\text{{F}LOPs}\rightarrow\text{{L}oss}$: Predict the target pre-training loss based on the expended FLOPs. Following previous work, we train a series of sampling LMs within the same model family to develop a power-law predictive model. For this stage, the expended FLOPs are not required to reach above the emergent threshold. (2) $\text{{L}oss}\rightarrow\text{{P}erformance}$: Predict the downstream performance based on the pre-training loss. We collect data points from intermediate checkpoints of various sampling LMs that exhibit above-random performance, and develop a regression model for prediction. In preliminary experiments with sampling LMs up to 3B, this FLP solution predicts the performance of 7B and 13B LLMs across various benchmarks with error margins of 5% and 10% respectively, significantly outperforming direct FLOPs-to-Performance predictions.
@@
Motivated by these findings, we present FLP-M, a fundamental solution for performance prediction that addresses the growing demand for integrating diverse datasets during LLMs pre-training, focusing on integrating the general corpus with code data in this work. FLP-M targets fine-grained domain-specific pre-training loss to capture the performance changes. Specifically, we extend the power law analytical function to predict the domain-specific loss based on FLOPs across multiple data sources. Then we employ a two-layer neural network to model the non-linear relationship between multiple domain-specific loss and the downstream performance. Through evaluation, we demonstrate that FLP-M effectively predicts the performance of 3B and 7B LLMs trained on various data mixtures (within 10% error margins for most benchmarks). This is achieved by utilizing a 3B LLM trained on a specific data mixing ratio along with a series of smaller sampling LMs.
2 Related Work
2.1 Scaling Laws
2.2 Data Mixture
@@ 3 FLP: Downstream Performance Prediction
We introduce a two-stage approach to predicting downstream performance in LLMs based on two established findings: (1) Predicting the target pre-training loss and establishing the power-law relation is feasible as it does not involve an emergent phase (Kaplan et al., 2020; Hoffmann et al., 2022). (2) When pre-training loss goes below a task-specific threshold, there is an observed correlation between pre-training loss and downstream task performance (Du et al., 2024; Huang et al., 2024). In this section, we present FLP as a proof-of-concept for this framework with a straightforward implementation.
@@ 3.1 FLOPs $\rightarrow$ Loss
where $C_{N}$ and $\alpha_{N}$ are constant terms to be estimated. In FLP, we train a series of $N$ LMs within the same model family in the same pre-training distribution, progressively increasing model size and training tokens to achieve even sampling. Then we measure their pre-training loss in our curated validation dataset to obtain $N$ pairs of ($C_{i}$, $L_{i}$) to estimate the constants in Eq. 1.
@@ 3.2 Loss $\rightarrow$ Performance
Based on our empirical observation of the scatter plots showing (pre-training loss, performance) data points (see §A), we select the analytical linear function to characterize the relation between the pre-training loss $L$ on general validation data and the task performance $P$:
@@ 4.1 Sampling and Target LMs
We train a series of 12 sampling LMs up to 3B parameters to predict the performance of target LLMs with 7B and 13B parameters. The configurations of LMs are shown in Tab. 1. We first determine the number of training tokens required for the 7B LLM (approximately 180 times the model size), considering practical needs and inference-time costs. In real-world applications, prioritizing inference efficiency often involves training smaller LMs with a higher token-to-parameter ratio beyond the optimal factor of 20x (Hoffmann et al., 2022). Our preliminary experiments indicate that scaling laws remain applicable even in this over-training regime (within 2.8% error margins). We then proportionally scale down this number to determine the required training tokens for the sampling LMs.
@@ 4.3 Experimental Setting
where $C_{M}$ and $\alpha_{M}$ are constant terms to be estimated. We denote this approach as FP.
@@ 4.4 Results
In addition, we discuss additional results in Appendix for the presentation purpose since adding these data points may distort the vertical axis scaling in Fig. 2. We compare FLP further with the analytical forms and approaches proposed in GPT-4 (Achiam et al., 2023) and Llama-3 (Dubey et al., 2024) technical reports. The results are shown in §B and §C respectively. We also evaluate the feasibility of employing FLP to predict the performance of a 13B LLM on MMLU (Hendrycks et al., 2020), using intermediate checkpoints from a 7B LLM (§D). Overall, the results demonstrate the general effectiveness and applicability of FLP.
@@ 5 FLP-M: Data Mixing for Downstream Performance Prediction
Motivated by the encouraging results of FLP4), we propose FLP-M, a fundamental approach to meet the practical needs of integrating data from various sources (Groeneveld et al., 2024; Penedo et al., 2024). In our work, we focus on mixing general corpus with code data, considering two distinct yet overlapping data sources. This intersection offers a more realistic perspective than treating them as distinct domains (Ye et al., 2024), as real-world corpus often spans multiple domains, necessitating an analysis of the interdependence between data sources when formulating our analytical functions.
Compared to the straightforward implementation of FLP3), FLP-M operates on fine-grained, domain-specific pre-training loss, due to the observation that the average loss on the entire validation set fails to effectively reflect performance variations in downstream tasks in the data mixing context (§7.2). This may be due to the fact that changes in pre-training data mixtures simultaneously impact multiple capabilities of the LMs. For instance, an increase in code data loss coupled with a decrease in general data loss may leave the average validation loss unchanged, yet result in LMs with distinct capabilities and downstream performance. Note that unlike the pre-training data mixture, the validation set is deliberately curated by domain, as creating smaller, domain-specific validation sets is manageable.
@@ 7.2 Ablation Study
FLP-M: FLOPs $\rightarrow$ Loss We experiment with several candidate analytical forms listed in Tab. 3. We assess their performance in estimating the average pre-training loss across the entire validation set, as well as the domain-specific pre-training losses on corresponding subsets. We present the fit curves in Fig. 13E), and the relative prediction errors for pre-training loss estimation are shown in Fig. 7. For average pre-training loss prediction, using more complex analytical models that account for the individual impact of each data source can lead to performance degradation. However, relying solely on the total compute for prediction (M1) can cause high prediction errors in certain domains (e.g., code) and are not stable for various mixing ratios. More complex analytical models generally perform better in predicting domain-specific loss. Among them, M4, the adopted model in FLP-M, provides more stable (within 2.5% relative prediction error across most domains) and overall more accurate predictions (achieving the lowest average error shown in Tab. 3).
@@ 8 Conclusion
This paper introduces a two-stage FLP solution to predict downstream performance in LLMs by leveraging pre-training loss. Encouraged by promising preliminary results, we propose FLP-M, a core solution for performance prediction that addresses the practical challenges of integrating pre-training data from diverse sources. The effectiveness of FLP-M is validated through extensive experiments.
@@
Our approach FLP-M is generally applicable across various data sources, yet currently, it is demonstrated only in binary cases involving code and text data due to computational constraints. Our specific emphasis on the mixing ratio of code is deliberate, reflecting its practical significance in real-world applications. This limitation marks a key area for future expansion.
@@
Appendix A Linear Relation Between Loss and Performance
We gather data points from intermediate checkpoints of all sampling LMs and visualize the relationship between pre-training loss and corresponding task performance in Fig. 9. We observe a generally linear trend across all benchmarks, which motivates our selection of linear analytical form to characterize the mapping from pre-training loss to downstream performance.
@@ Appendix A Linear Relation Between Loss and Performance
Appendix B Analytical Form to Fit FLOPs-to-Performance Curve

### v2

@@ Abstract
Precise estimation of downstream performance in large language models (LLMs) prior to training is essential for guiding their development process. Scaling laws analysis utilizes the statistics of a series of significantly smaller sampling language models (LMs) to predict the performance of the target LLM. For downstream performance prediction, the critical challenge lies in the emergent abilities in LLMs that occur beyond task-specific computational thresholds. In this work, we focus on the pre-training loss as a more computation-efficient metric for performance estimation. Our two-stage approach FLP consists of first estimating a function that maps computational resources (e.g., FLOPs) to the pre-training Loss using a series of fully-converged sampling models, followed by mapping the pre-training loss to downstream task Performance using the intermediate models with emerged performance. In our experiments, this FLP solution accurately predicts the performance of LLMs with 7B and 13B parameters using a series of sampling LMs up to 3B, achieving error margins of 5% and 10%, respectively, and significantly outperforming the FLOPs-to-Performance approach. Further, we present FLP-M, a fundamental approach for performance prediction that addresses the practical need to integrate datasets from multiple sources during pre-training, specifically blending general corpus with code data to accurately represent the common necessity. FLP-M extends the power law analytical function to predict domain-specific pre-training loss based on FLOPs across data sources, and employs a two-layer neural network to model the non-linear relationship between multiple domain-specific loss and downstream performance. By utilizing a 3B LLM trained on a specific ratio and a series of smaller sampling LMs, FLP-M can effectively forecast the performance of 3B and 7B LLMs across various data mixtures for most benchmarks within 10% error margins.
@@ 1 Introduction
Previous efforts focus on predicting the target LLM’s pre-training loss and establish a power-law relation between the computational resource expended (e.g., floating-point operations per second (FLOPs)) and the final loss achieved (Kaplan et al., 2020; Muennighoff et al., 2024; Henighan et al., 2020). Further, we aim to predict the downstream performance in pre-trained LLMs (i.e., zero- or few-shot evaluation) to more accurately reflect the primary concerns regarding their capabilities. The critical challenge is the emergent abilities in LLMs, which states that LLMs only exceed random performance when the FLOPs expended during training surpass task-specific thresholds (Wei et al., 2022a).
@@
Supposing a task threshold of $F_{c}$, typical methods require training $N$ LMs, expending total FLOPs $F_{t}=\sum_{i=1}^{N}\text{FLOPs}_{i}>N\times F_{c}$, to obtain $N$ effective data points, thereby necessitating significant computational resources. Fig. 1 demonstrates that the sampling LMs require more than $5\times 10^{20}$ FLOPs to perform better than random on most benchmarks, with only three data points available to fit the predictive curve across these benchmarks. Hu et al. (2023) address this challenge by significantly increasing the sampling times to compute the PassUntil of a task, basically increasing the “metric resolution” to enable the abilities to emerge earlier (i.e., reducing $F_{c}$). However, this approach faces challenges in translating the PassUntil back to the original task metric of concerns and requires huge amounts of FLOPs spent on sampling.
@@
In this work, we target the actual task performance prediction based on two intuitions: (1) Predicting the target pre-training loss is easier and achievable since there is no “emergent phase” in the pre-training loss, as extensively verified in Kaplan et al. (2020); Hoffmann et al. (2022); (2) There is an observed correlation between the pre-training loss and the downstream task performance after the “emergent point” (i.e., the pre-training loss goes below a critical threshold) (Du et al., 2024; Huang et al., 2024). Specifically, LMs that achieve comparable pre-training loss demonstrate consistent performance levels on downstream tasks, regardless of variations in hyper-parameter configurations (e.g., model size, the number of training tokens, and learning rate schedules). Thus, different from estimating the FLOPs-to-Loss function that requires training $N$ LMs to convergence, we can collect (pre-training loss, performance) data points at intermediate checkpoints that have emerged performance to estimate the Loss-to-Performance function, thereby enhancing sample efficiency.
@@
Formally, our approach, named FLP, consists of two sequential stages: (1) $\text{{F}LOPs}\rightarrow\text{{L}oss}$: Predict the target pre-training loss based on the expended FLOPs. Following previous work, we train a series of sampling LMs within the same model family and collect data points from their final fully-converged checkpoints to develop a power-law predictive model. For this stage, the expended FLOPs are not required to reach above the task-specific emergent threshold. (2) $\text{{L}oss}\rightarrow\text{{P}erformance}$: Predict the downstream performance based on the pre-training loss. We collect data points from intermediate checkpoints of various sampling LMs that exhibit above-random performance, and develop a regression model for prediction. In our experiments with sampling LMs up to 3B, this FLP solution predicts the performance of 7B and 13B LLMs across various benchmarks with error margins of 5% and 10% respectively, significantly outperforming direct FLOPs-to-Performance predictions.
@@
Motivated by these findings, we present FLP-M, a fundamental solution for performance prediction that addresses the growing demand for integrating diverse datasets during LLMs pre-training, focusing on integrating the general corpus with code data in this work. FLP-M targets fine-grained domain-specific pre-training loss to capture the performance changes. Specifically, we extend the power law analytical function to predict the domain-specific loss based on FLOPs across multiple data sources. Then we employ a two-layer neural network to model the non-linear relationship between multiple domain-specific loss and the downstream performance. Through evaluation, we demonstrate that FLP-M effectively predicts the performance of 3B and 7B LLMs trained on various data mixtures (within 10% error margins for most benchmarks). This is achieved by utilizing a 3B LLM trained on a specific data mixing ratio along with a series of smaller sampling LMs. Furthermore, our extensive experiments demonstrate that FLP-M effectively optimizes data mixture compositions to improve downstream task performance. Through comprehensive ablation studies, we validate the design choices in our analytical functions.
Overall, this work presents three major contributions to LLMs development. First, we introduce a systematic and principled framework for forecasting downstream performance in LLMs. Second, we innovatively employ pre-training loss as an intermediate variable, enabling both the handling of training dynamics and quantitative monitoring of downstream performance. Third, we present a method extension to manage data mixing scenarios, which is crucial for LLMs pre-training.
@@ 2 FLP: Downstream Performance Prediction
We introduce FLP, a two-stage approach to predicting downstream performance in LLMs based on two established findings: (1) Predicting the target pre-training loss and establishing the power-law relation is feasible as it does not involve an emergent phase (Kaplan et al., 2020; Hoffmann et al., 2022). (2) When pre-training loss goes below a task-specific threshold, there is an observed correlation between pre-training loss and downstream task performance (Du et al., 2024; Huang et al., 2024) regardless of the hyper-parameter configurations.
@@ 2.1 FLOPs $\rightarrow$ Loss
where $C_{N}$ and $\alpha_{N}$ are constant terms to be estimated. In FLP, we train a series of $N$ LMs within the same model family in the same pre-training distribution, progressively increasing model size and training tokens to achieve even sampling. Then we measure their pre-training loss in our curated validation dataset to obtain $N$ pairs of ($C_{i}$, $L_{i}$) to estimate the constants in Eq. 1. Note that the data points are only collected from the final fully-converged sampling LMs, which ensures the one-to-one mapping from FLOPs to pre-training loss (Kaplan et al., 2020).
@@ 2.2 Loss $\rightarrow$ Performance
Based on our empirical observation of the scatter plots showing (pre-training loss, performance) data points (see §A) and the strong statistical evidence that the average coefficient of determination $R^{2}$ is 93% across all benchmarks for the linear fitting curves, we select the analytical linear function [^1] In our preliminary investigations, we also explore nonlinear alternatives, including the sigmoid function, though these approaches demonstrate suboptimal predictive performance. to characterize the relation between the pre-training loss $L$ on general validation data and the task performance $P$:
@@
2.3 The Necessity of Two-Stage Modeling
A natural question is why we cannot simply combine Eq. 1 and Eq. 2 to directly model the relationship between FLOPs and performance. The key issue is that the relationship between FLOPs and task performance is not a simple direct mapping. For a given FLOPs budget, the resultant task performance can vary significantly depending on the training dynamics (e.g., learning rate schedule, model size, and the number of training tokens). This many-to-one relationship between FLOPs and performance makes fitting a direct analytical function between them unreliable. Our two-stage approach addresses this by using pre-training loss as an intermediate variable. Critically, we can reliably fit the Loss-to-Performance mapping using (pre-training loss, performance) data points, since models with similar pre-training loss tend to demonstrate comparable task performance, regardless of their specific training trajectories. However, we cannot fit the direct FLOPs-to-performance mapping using (FLOPs, performance) data points due to the variation in training dynamics. Overall, the two-stage formulation allows us to more accurately model and optimize the relationship between computational resources and model performance. We implement a baseline in §3 for comparison to demonstrate that directly composing Eq. 1 and Eq. 2 cannot fully capture the complex training dynamics accurately.
@@ 3.1 Sampling and Target LMs
We train a series of 12 sampling LMs up to 3B parameters to predict the performance of target LLMs with 7B and 13B parameters. The configurations of LMs are shown in Tab. 1. We adopt the fixed data-model ratio scaling strategy (Kaplan et al., 2020; Hoffmann et al., 2022) that maintains a fixed ratio between training tokens and model size while varying compute (FLOPs). In this scaling strategy, for any given compute budget, there exists a pre-determined allocation between training tokens and model size. We first determine the number of training tokens required for the 7B LLM (approximately 180 times the model size), considering practical needs and inference-time costs. In real-world applications, prioritizing inference efficiency often involves training smaller LMs with a higher token-to-parameter ratio beyond the optimal factor of 20x (Hoffmann et al., 2022). Our preliminary experiments indicate that scaling laws remain applicable even in this over-training regime (within 2.8% error margins). We then proportionally scale down this number to determine the required training tokens for the sampling LMs.
@@ 3.3 Experimental Setting
where $C_{M}$ and $\alpha_{M}$ are constant terms to be estimated. We denote this approach as FP. For a fair comparison, we collect data points using identical sampling LMs across both FLP and FP.
@@ 3.4 Results
In addition, we discuss additional results in Appendix for the presentation purpose since adding these data points may distort the vertical axis scaling in Fig. 2. We compare FLP further with the analytical forms and approaches proposed in GPT-4 (Achiam et al., 2023) and Llama-3 (Dubey et al., 2024) technical reports. The results are shown in §C and §D respectively. We also evaluate the feasibility of employing FLP to predict the performance of a 13B LLM on MMLU (Hendrycks et al., 2020), using intermediate checkpoints from a 7B LLM (§B). Overall, the results demonstrate the general effectiveness and applicability of FLP.
@@ 4 FLP-M: Data Mixing for Downstream Performance Prediction
Motivated by the encouraging results of FLP3), we propose FLP-M, a fundamental approach to meet the practical needs of integrating data from various sources (Groeneveld et al., 2024; Penedo et al., 2024). In our work, we focus on mixing general corpus with code data, considering two distinct yet overlapping data sources. This intersection offers a more realistic perspective than treating them as distinct domains (Ye et al., 2024), as real-world corpus often spans multiple domains, necessitating an analysis of the interdependence between data sources when formulating our analytical functions.
Compared to the straightforward implementation of FLP2), FLP-M operates on fine-grained, domain-specific pre-training loss, due to the observation that the average loss on the entire validation set fails to effectively reflect performance variations in downstream tasks in the data mixing context (§6.2). This may be due to the fact that changes in pre-training data mixtures simultaneously impact multiple capabilities of the LMs. For instance, an increase in code data loss coupled with a decrease in general data loss may leave the average validation loss unchanged, yet result in LMs with distinct capabilities and downstream performance. Note that unlike the pre-training data mixture, the validation set is deliberately curated by domain, as creating smaller, domain-specific validation sets is manageable.
@@ 4.2 Domain Loss $\rightarrow$ Performance
Note that while neural networks are well-suited for the stage-2 Loss-to-Performance mapping due to abundant data points collected from intermediate checkpoints, they are less appropriate for stage-1 FLOPs-to-Loss mapping because it requires data points from fully-converged models, yielding limited data points from our configurations (Tab. 1).
@@ 6.2 Ablation Study
FLP-M: FLOPs $\rightarrow$ Loss We experiment with several candidate analytical forms listed in Tab. 3. We assess their performance in estimating the average pre-training loss across the entire validation set, as well as the domain-specific pre-training losses on corresponding subsets. We present the fit curves in Fig. 14E), and the relative prediction errors for pre-training loss estimation are shown in Fig. 7.
@@
For average pre-training loss prediction, using more complex analytical models that account for the individual impact of each data source can lead to performance degradation. However, relying solely on the total compute for prediction (M1) can cause high prediction errors in certain domains (e.g., code) and are not stable for various mixing ratios. More complex analytical models generally perform better in predicting domain-specific loss. Among them, M4, the adopted model in FLP-M, provides more stable (within 2.5% relative prediction error across most domains) and overall more accurate predictions (achieving the lowest average error shown in Tab. 3).
@@
6.3 Using Domain Loss in FLP
We explore the application of FLP-M when pre-training on a consistent distribution (the experimental setting described in §3), and compare it with FLP. The fitting curves are shown in Fig. 16 (Appendix) and the results of relative prediction error are shown in Fig. 9. We show that FLP-M fails to effectively predict the performance of target LLMs when sampling LMs are pre-trained on a fixed distribution. This ineffectiveness is attributed to the closely related domain-specific validation losses among the sampling LMs within the same training distribution, which suggests that decomposing the pre-training validation loss yields no additional information in this pre-training setting. Thus, estimating five domain-specific loss, rather than a single average validation loss, can further increase the risk of error propagation. Moreover, using highly correlated features as neural network inputs may lead to overfitting.
7 Related Work
7.1 Scaling Laws
7.2 Data Mixture
@@ 8 Conclusion
This paper introduces a two-stage FLP solution to predict downstream performance in LLMs by leveraging the pre-training loss to address the challenges of LLMs’ emergent abilities. In addition, we propose FLP-M, a core solution for performance prediction that addresses the practical challenges of integrating pre-training data from diverse sources. The effectiveness of FLP and FLP-M is validated through extensive experiments. Furthermore, we demonstrate both the effectiveness of our selected analytical forms and the utility of FLP-M in optimizing data mixtures to enhance downstream task performance.
@@
Our approach FLP-M is generally applicable across various data sources, yet currently, it is demonstrated only in binary cases involving code and text data due to computational constraints. Our specific emphasis on the mixing ratio of code is deliberate, reflecting its practical significance in real-world applications and the dominance of these two data sources in LLMs pre-training compared to others (e.g., academic articles, books, encyclopedias).
@@
Appendix A Linear Relation Between Loss and Performance
We gather data points from intermediate checkpoints of all sampling LMs and visualize the relationship between pre-training loss and corresponding task performance in Fig. 10. We observe a generally linear trend across all benchmarks, which motivates our selection of linear analytical form to characterize the mapping from pre-training loss to downstream performance. In addition, the decision is further supported by the following evidence:
  • Huang et al. (2024) demonstrate through extensive experiments across 12 benchmarks and 31 LLMs that language modeling ability (compression) correlates linearly with benchmark performance (intelligence).
  • Du et al. (2024) provide empirical validation and visualization of this linear relationship specifically in the post-emergence regime.
Overall, the above results and supporting evidence also justify that monitoring pre-training loss alone is sufficient to measure downstream benchmark performance in LLMs. Changes in pre-training loss correspond to proportional changes in downstream benchmark performance, following a consistent linear trend without significant variance.
Appendix B MMLU Experiment
Figure 13: The performance prediction on MMLU using FLP.
Appendix C Analytical Form to Fit FLOPs-to-Performance Curve